OpenStudy (paige):

6log9x + 3log27y=8 log3x+2log9y=2. Solve the equation ;)

myininaya (myininaya):

those numbers next to the variables aren't the bases right?

myininaya (myininaya):

the equations are exactly as you have them right?

OpenStudy (paige):

Yes. The numbers are the base.

myininaya (myininaya):

so we have \[6\log_9(x)+3\log_{27}(y)=8\] and \[\log_3(x)+2\log_9(y)=2\] ?

OpenStudy (paige):

Yes!

OpenStudy (anonymous):

Since all those bases are powers of 3, lets convert all the bases to 3 using this formula: \[\log_{b^{n}}x= \frac{1}{n}\log_{b}x\] 9 is 3^2, and 27 is 3^3, so after converting we get: \[\frac{6}{2}\log_{3}x+\frac{3}{3}\log_3y = 8\] \[\log_3x+\frac{2}{2}\log_3y=2\]

OpenStudy (anonymous):

now we have a system we can solve: \[3\log_3x+\log_3y=8\] \[\log_3x+\log_3y=2\] Subtracting the second equation from the first gets rid of the y part: \[2\log_3x=6 \Rightarrow \log_3x=3 \Rightarrow x = 27\] Then, if: \[\log_3x+\log_3y= 2 \Rightarrow 3+\log_3y = 2 \Rightarrow \log_3y = -1\] So y = 1/3

OpenStudy (paige):

When you get log3x=3, how do you find x?

myininaya (myininaya):

lol joe i made my more complicated then it needed to be but i did get what you did

OpenStudy (anonymous):

When looking at log functions this is how they should be interpreted: \[\log_b x= a \Leftrightarrow b^{a} = x\] So when I see: \[\log_3 x = 3\] I change it to: \[3^{3} = x \Rightarrow x = 27\]

OpenStudy (anonymous):

@myininaya yeah i was scribbling on paper and it was getting ridiculous >.< haha

myininaya (myininaya):

only read this if you want to get confused

OpenStudy (paige):

Ok. Thx for your help! I have exams in 2 hrs xD & I got the answers! :D

myininaya (myininaya):

i forgot to use that one propery ln3^3=3ln3 so i wrote it as ln3+ln3+ln3 then i was like oh yeah thats just 3ln3 lol

myininaya (myininaya):

but still even if i realized that you did something much better

OpenStudy (anonymous):

lol, I had started mine out the exact same way, and it just seemed too long, so i was grasping for straws trying to find a way to make the problem easier >.<